Solve a, b and 2

1. Calculate the pH during the titration of 10.00 mL of 0.400 M hypochlorous acid with 0.500 M NaOH. First what is the initial pH (before any NaOH is added)?
The Ka for HOCl is 3.0 x 10-8 M. Answer is 3.96

a. What is the pH after 10.60 mL of NaOH are added?

b. What is the pH after 15.30 mL of NaOH are added?

2. Chloropropionic acid, ClCH2CH2COOH is a weak monoprotic acid with Ka = 7.94 x 10-5 M. Calculate the pH at the equivalence point in a titration 41.0 mL of 0.60 M chloropropionic acid with 0.5 M KOH.
Answer #1

3-0xi08 M, pka - (1) Ka = - -log C3x108) IAi = 7.52 Initial pt-(a lag c) 1 t(7.52-lag (o4) =(7.52-leg(o-4)) = 1.92 2- = 3.96